# Point2Primitive: CAD Reconstruction from Point Cloud by Direct Primitive Prediction

### v2

@@ 4.3 Extrusion Segmentation
Our method utilizes a point segmentation network to cluster points and rebuild the topology, as shown in Figure 2. The Point2Primitive is insensitive to the implementation of the point segmentation network. We choose the PointNet++ [25] as the point encoder and the decoder, which is detailed in Appendix 4.1.
@@ 4.4 Improved Transformer Decoder
The curve parameters are formulated in a center-prior format detailed in Section 3, so the curve parameters can be formulated as position encoding $PE_{i}$ to guide the attention. Given the curve parameters $\rho^{i}_{j}$ as the parameters of the $j\mbox{-}th$ curve in the $i\mbox{-}th$ sketch $S_{i}$. The corresponding positional queries $PE_{j}^{i}$ are generated by:
$PE_{j}^{i}=\text{MLP}(f_{pe}(\rho^{i}_{j}))=\text{MLP}(f_{pe}(x_{m_{j}}^{i},y_{m_{j}}^{i}))$ (3)
where $f_{pe}$ is the function that generates the sinusoidal embeddings from the curve center. Multi-layer perception (MLP) and ReLU activations are used to produce the position encodings $PE_{j}^{i}$.
@@
$\delta x_{m_{j}}^{i},\delta y_{m_{j}}^{i},\delta p_{2_{j}}^{i},\delta p_{3_{j}}^{i},\delta p_{4_{j}}^{i},\delta p_{5_{j}}^{i}=\text{MLP}(O_{n})$ (4)
where $O_{n}$ is the output of the $n\mbox{-}th$ decoder layer. The curve parameters are updated to $[x_{m_{j}}^{i^{\prime}},y_{m_{j}}^{i^{\prime}},p_{2_{j}}^{i^{\prime}},p_{3_{j}}^{i^{\prime}},p_{4_{j}}^{i^{\prime}},p_{5_{j}}^{i^{\prime}}]$ in each layer, as shown in Figure 3. All intermediate parameter outputs are supervised by ground truth. DeNoising [12] is utilized to accelerate the convergence.
4.5 Loss Function
The Point2Primitive is trained using a multi-task, non-convex objective composed of extrusion segmentation loss $\mathcal{L}_{\mathbb{E}}$, base-barrel classification loss $\mathcal{L}_{{bb}}$, normal $\mathcal{L}_{norm}$ and sketch $\mathcal{L}_{skh}$ prediction losses:
$\mathcal{L}=\mathcal{L}_{\mathbb{E}}+\mathcal{L}_{{bb}}+\mathcal{L}_{norm}+\mathcal{L}_{skh}$ (5)
where the cross entropy loss is used in $\mathcal{L}_{\mathbb{E}}$ and $\mathcal{L}_{{bb}}$ while L-1 loss is used in $\mathcal{L}_{norm}$. The sketch prediction loss $\mathcal{L}_{skh}$ can be formulated as:
$\mathcal{L}_{skh}$ $=\sum\limits_{i=1}^{N_{C}}l_{focal}(\hat{t}_{i},t_{i})+\beta\sum\limits_{i=1}^{N_{C}}\sum\limits_{j=1}^{N_{p}}\mathcal{L}_{p}(\hat{\rho}_{i,j},\rho_{i,j})$ (6)
$\mathcal{L}_{p}$ $=L_{1}(\hat{\rho}_{i,j},\rho_{i,j})+L_{2}(\hat{\rho}_{i,j},\rho_{i,j})$
where $l_{focal}(\cdot,\cdot)$ represents the focal loss [20] and $L_{1}(\cdot,\cdot)$ and $L_{2}(\cdot,\cdot)$ represents the L1 and L2 loss, respectively. Both Denoising Loss and Curve Loss use the sketch prediction loss $\mathcal{L}_{skh}$, as shown in Figure 2. More specifically, in the L2 loss, we randomly select $n$ points on each curve and compute the chamfer distance as the L2 loss. $N_{P}$ is the number of parameters ($N_{p}=6$ is this paper) while $N_{C}=30$ is the number of curves. $\beta$ is the weight to balance both terms ($\beta=2$ in this paper) and $\hat{\cdot}$ denotes the ground truth value.
@@ 5.1 Experimental Setups
Datasets. To evaluate the proposed model, we conduct experiments on DeepCAD [37] and Fusion 360 Gallery [36]. The training data is generated as described in Sec. 4.2.
Implementation Details. We use PointNet++[25] for extrusion segmentation. The Point2Primitive is trained for 400 epochs with a total batch size of 64 and learning rate (lr) $51e-$ 5 with linear warmup and step reduce lr scheduler. As for the DeNoising settings, the noise rate for the curve type and parameters is $0.2$ and $0.3$, respectively, with 5 DeNoising groups. The transformer encoder and decoder contain 6 and 8 transformer layers, respectively, with eight attention heads and latent model dimension 256.
Evaluation Metrics. We adopt command type accuracy ($Acc^{SKH}_{t}$) and parameter accuracy ($Acc^{SKH}_{p}$) for quantitative evaluations of the sketch prediction following [5]. For networks that represent sketch by the implicit nueral field, the tools for convert sketch SDF to command are developed based on the tool by SECAD [14]. Furthermore, following [3], we also report chamfer distance ($CD$), edge chamfer distance ($ECD$), normal consistency ($NC$), and the number of generated primitives ($\Delta\#P$) to measure the quality of the recovered 3D geometry, more details of the quantitative metrics are shown in the Appendix 1.
@@ 5.2 Main Results
[b] DeepCAD [37] Fusion 360 Gallery [36] Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$ IR #$\#\Delta P$ $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$ IR #$\#\Delta P$ Methods of Representing Sketch with Implicit SDF Encoding ExtrudeNet [29] 34.58% 31.71% 0.379 0.962 0.849 24.19% 34.34 37.81% 34.39% 0.671 0.808 0.809 23.83% 27.47 Point2Cyl [34] 41.37% 39.41% 0.489 1.027 0.819 3.91% 26.18 42.98% 41.16% 0.529 0.769 0.769 3.87% 26.96 SECAD-Net [14] 45.91% 41.37% 0.341 0.868 0.861 8.03% 32.78 46.82% 42.97% 0.449 0.684 0.813 7.82% 28.61 Generation Methods Based on Language Model DeepCAD [37] 82.61% 73.36% 0.898 1.883 0.823 14.12% 5.89 77.31% 69.81% 7.128 8.729 0.719 13.18% 7.47 HNC-CAD [39] 84.31% 76.71% 0.827 1.064 0.846 6.01% 4.43 80.62% 72.19% 4.381 5.571 0.748 5.92% 6.14 Methods Based on Primitive Fitting UCSG-Net [7] - - 1.849 1.174 0.820 - 14.19 - - 0.952 1.277 0.770 - 11.17 CSG-Stump [28] - - 3.031 0.755 0.828 - 19.46 - - 0.781 0.991 0.744 - 13.08 Our Method 96.14% 86.81% 0.312 0.581 0.897 3.71% 4.14 94.17% 83.52% 0.392 0.571 0.839 3.62% 5.17
@@
To demonstrate its effectiveness, we compare the proposed model to multiple existing methods, covering the methods of representing sketch with SDF encoding [14, 29, 34], the generation methods based on language model [37, 39], and methods based on primitive fitting [7, 28]. The direct reconstruction results are visualized for comparison.
Table 2 summarizes the main results, and we can find that sketch curve type and parameter accuracy of the Point2Primitive are much higher than the other methods while the CD metric of the Point2Primitive is lower. The smallest CD and ECD value indicates that the presented CAD reconstruction method achieves better geometry fidelity while preserving accurate and sharp shape edges. This shows that by directly recovering curves of the extrusion primitive, our method can achieve better performance than the other extrusion-segmentation methods that represent the sketch with an implicit field. As for the number of generated primitives, our method is closer to human designs (The smallest #$\#\Delta P$ value). This improvement is due to the explicit fine-grained reconstruction from each extrusion primitive; every parameter of the curves and extrusion operations are all predicted and optimized.
@@ 5.3 Visualization Results
Figure 4: Visual comparison between reconstruction results on the DeepCAD (left) and Fusion360 gallery dataset (right).
Figure 4 demonstrates the visualized reconstruction results on the DeepCAD and Fusion 360 Gallery datasets, respectively. It can be seen that the results of our method are more compact and complete, while the edges are sharp and sketches are accurate. This shows that our method can produce CAD reconstruction of high geometrical fidelity. The methods based on the language model can also produce accurate geometry structure, but the error of the parameters in the sequence results in some models of poor compactness. The primitive detection methods (USCG and Stump) fail to predict some holes in the model. Still, the edges are critically more accurate than the ExtrudeNet, whose reconstructions are less complete.
@@ 5.4 Comparison on Augmented DeepCAD Dataset
Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ $CD$ $ECD$ $NC$$\#\Delta P$
Methods of Representing Sketch with Implicit SDF Encoding
ExtrudeNet [29] 28.39% 25.61% 0.614 1.117 0.776 36.14
Point2Cyl [34] 34.47% 30.25% 0.518 1.065 0.791 27.96
SECAD-Net [14] 36.61% 31.43% 0.437 1.079 0.806 34.18
Generation Method Based on Language Model
DeepCAD [37] 61.41% 43.71% 5.919 6.883 0.708 7.16
HNC-CAD [39] 63.39% 53.29% 6.864 7.064 0.711 6.56
Methods Based on Primitive Fitting
UCSG-Net [7] - - 2.146 1.273 0.797 14.81
CSG-Stump [28] - - 3.681 0.958 0.804 20.16
Our Method 94.67% 83.81% 0.410 0.607 0.819 4.97
Table 3: Evaluation results on the Augmented DeepCAD dataset
@@
Table 3 demonstrates the quantitative results of the augmented DeepCAD dataset. We add the original shape (Origin) to the last column in the visualized results as shown in Figure 5. It can be seen that the methods based on the language model see a direct drop in the metrics of the sketch command type and command parameter, leading to a significant increase in the CD metric. Some reconstruction results by the language model preserve similarities to the original shape, leading to false reconstruction. The generated results are similar to the ground truth but with low geometry accuracies. On the other hand, reconstruction methods can still produce results that satisfy geometry fidelity. This implies that the methods based on the language model are less generalizable than the other methods. Also, both the quantitative metrics and the visual results verify that the proposed method can be generalized to reconstruction tasks beyond the train-test dataset.
@@ 5.5 Detailed Comparison with Point2Cyl
Figure 6: Comparision with the Point2Cyl: (a) The network architecture of Point2Cyl (left) and our Point2Primitive (right); (b) The reconstructed sketch by Point2Primitive and Point2Cyl; (c) The extrusion segmentation by Point2Primitive and Point2Cyl.
Point2Cyl is the most related method to the proposed Point2Primitive, so we conduct extra comparisons with Point2Cyl. Figure 6 demonstrates the architecture difference between Point2Cyl and the proposed Point2Primitive. Instead of using IGR [6] and PointNet[2] to encode SDF for sketch SDF prediction, an improved transformer is utilized in the Point2Primtive to convert barrel points to curves directly. Thus, the sketch loss is the curve prediction loss rather than the regression loss of the sketch SDF encodings, which introduces more explicit geometry information. Also, we utilize the technique based on the jet fitting [1] (for sketch prediction) and Hough Transform [19] (for extrusion axis) as the optimization methods for comparison.
Metrics Ours Ours Point2Cyl[34] Point2Cyl[34] JF[1]
w/o $\mathcal{L}_{sketch}$ w/o $\mathcal{L}_{sketch}$ HT[19]
Ext. Seg IoU 0.814 0.862 0.801 0.817 0.513
BB. Seg Acc 0.902 0.902 0.867 0.867 0.601
EA. Angle Error 7.416 7.173 8.391 8.267 57.147
Fit Ext. 0.0771 0.0527 0.0791 0.0741 0.1701
Table 4: Comparison with Point2Cyl on DeepCAD dataset.
Table 4 demonstrates the quantitative comparison. Compared with Point2Cyl, the metrics of the extrusion (Ext.) and barrel/based (BB.) segmentation are improved by $5.6\%$ and $3.5\%$, respectively, leading to $13\%$ and $32\%$ reduction on the extrusion axis (EA.) angle and fitting (Fit Ext.) error, respectively. This shows that by supervising the extrusion segmentation with curve loss (type and parameter loss), more geometrical information is introduced, leading to better geometry accuracy. We visualized the sketch predicted by Point2Cyl and Point2Primitive, as shown in Figure 6 (b). It can be seen that because Point2Cyl represents the sketch by SDF, the edges of the sketch profiles are curved, while the counterpart of Point2Primitive is sharp and accurate. Further, we present more visualizations of the predicted sketch by Point2Primitive in the Appendix 2. Figure 6 (c) demonstrates the segmentation results by the two methods. It can be seen that even though the multiple parts of one extrusion are labeled with different extrusion labels (rows 4 and 5), the segmentation network can still learn the geometry representation and cluster the points as the same extrusion. This implies that Point2Primitive can predict simplified primitives that are closer to human design.

### v3

@@ 3.3 Extrusion Segmentation
To decompose the raw point cloud into semantically meaningful extrusion bodies, we design an extrusion segmentation network that clusters points belonging to the same extrusion. As illustrated in Figure 1, the segmentation module is implemented using a PointNet++ encoder–decoder backbone (Qi_Yi_Su_Guibas_2017), followed by two prediction heads.
Specifically, the encoder extracts hierarchical features from the input point cloud, while the decoder progressively recovers point-wise embeddings. On top of these embeddings, two parallel branches are employed: the Normal Head predicts point-wise surface normals, and the Extrusion Segmentation Head predicts extrusion-related classes. The Ext Head outputs logits $\hat{\mathbf{M}}$ corresponding to $2K$ categories, where each extrusion is represented by two complementary channels. For the $j$-th extrusion, the $(2j)$-th channel encodes the extrusion body assignment, whereas the $(2j+1)$-th channel encodes its associated barrel points. In this way, extrusion logits $\hat{\mathbf{W}}$ (for generating the extrusion $\textbf{p}_{k}$) and barrel logits $\hat{\mathbf{B}}$ (for defining the the boundary $b$) can be computed as:
$\hat{\mathbf{W}}_{:,j}$ $=\hat{\mathbf{M}}_{:,2j}+\hat{\mathbf{M}}_{:,2j+1},$ (3)
$\hat{\mathbf{B}}_{:,0}$ $=\sum_{j}\hat{\mathbf{M}}_{:,2j},\quad\text{and}\quad\hat{\mathbf{B}}_{:,1}=\sum_{j}\hat{\mathbf{M}}_{:,2j+1}.$ (4)
Through this design, the network jointly captures extrusion topology and barrel boundary information, which facilitates accurate decomposition of the complex shape into extrusion units. Please refer to Appendix E for more implementation detail.
@@ 3.4 Improved Transformer Decoder
The curve parameters are formulated in a center-prior format detailed in Section 3.1, so the curve parameters can be formulated as position encoding $\textbf{PE}_{i}$ to guide the attention. Given the curve parameters $\rho^{i}_{j}$ as the parameters of the $j\mbox{-}th$ curve in the $i\mbox{-}th$ sketch $S_{i}$. The corresponding positional queries $\textbf{PE}_{j}^{i}$ are generated by:
$\textbf{PE}_{j}^{i}=\texttt{MLP}(f_{\text{pe}}(\rho^{i}_{j}))=\texttt{MLP}(f_{\text{pe}}(x_{m_{j}}^{i},y_{m_{j}}^{i})),$ (5)
where $f_{pe}$ is the function that generates the sinusoidal embeddings from the curve center. Multi-layer perception (MLP) and ReLU activations are used to produce the position encodings $\textbf{PE}_{j}^{i}$.
@@
$\delta x_{m_{j}}^{i},\delta y_{m_{j}}^{i},\delta p_{2_{j}}^{i},\delta p_{3_{j}}^{i},\delta p_{4_{j}}^{i},\delta p_{5_{j}}^{i}=\texttt{MLP}(O_{n}),$ (6)
where $O_{n}$ is the output of the $n\mbox{-}th$ decoder layer. The curve parameters are updated to $x_{m_{j}}^{i^{\prime}},y_{m_{j}}^{i^{\prime}},p_{2_{j}}^{i^{\prime}},p_{3_{j}}^{i^{\prime}},p_{4_{j}}^{i^{\prime}},p_{5_{j}}^{i^{\prime}}]$ in each layer, as shown in Figure 2. All intermediate parameter outputs are supervised by ground truth. We further utilize query denoising (Li_Zhang_Liu_Guo_Ni_Zhang) to accelerate the convergence.
3.5 Training Objective
The Point2Primitive  is trained using a multi-task objective composed of extrusion segmentation loss $\mathcal{L}_{\text{ext}}$, base-barrel classification loss $\mathcal{L}_{\text{bb}}$, normal $\mathcal{L}_{\text{norm}}$ and sketch $\mathcal{L}_{\text{skh}}$ prediction losses:
$\mathcal{L}=\mathcal{L}_{\text{ext}}+\mathcal{L}_{\text{bb}}+\mathcal{L}_{\text{norm}}+\mathcal{L}_{\text{skh}},$ (7)
where the cross entropy loss is used in $\mathcal{L}_{\text{ext}}$ and $\mathcal{L}_{\text{bb}}$ while L-1 loss is used in $\mathcal{L}_{\text{norm}}$. The sketch prediction loss $\mathcal{L}_{\text{skh}}$ can be formulated as:
$\mathcal{L}_{skh}$ $=\sum\limits_{i=1}^{N_{C}}\mathcal{L}_{\text{focal}}(\hat{t}_{i},t_{i})+\beta\sum\limits_{i=1}^{N_{C}}\sum\limits_{j=1}^{N_{p}}\mathcal{L}_{\text{p}}(\hat{\rho}_{i,j},\rho_{i,j}),$ (8)
$\mathcal{L}_{p}$ $=\mathcal{L}_{1}(\hat{\rho}_{i,j},\rho_{i,j})+\mathcal{L}_{1}(\hat{\rho}_{i,j},\rho_{i,j}),$
where $\mathcal{L}_{\text{focal}}(\cdot,\cdot)$ represents the focal loss (Lin_Goyal_Girshick_He_Dollar_2017) and $\mathcal{L}_{1}(\cdot,\cdot)$ and $\mathcal{L}_{2}(\cdot,\cdot)$ represents the L1 loss and L2 loss, respectively. Both denoising loss and curve Loss use the sketch prediction loss $\mathcal{L}_{skh}$, as shown in Figure 1. More specifically, in the L2 loss, we randomly select $n$ points on each curve and compute the chamfer distance as the L2 loss. $N_{P}$ is the number of parameters ($N_{p}=6$ is this paper) while $N_{C}=30$ is the number of curves. $\beta$ is the weight to balance both terms ($\beta=2$ in this paper) and $\hat{\cdot}$ denotes the ground truth value.
@@ 4.1 Experimental Setups
Datasets. To evaluate the proposed model, we conduct experiments on DeepCAD (Wu_Xiao_Zheng_2021) and Fusion 360 Gallery (willis2021fusion360gallerydataset). The training data is generated as described in Sec. 3.2.
Implementation Details. We use PointNet++ (Qi_Yi_Su_Guibas_2017) for extrusion segmentation. The Point2Primitive  is trained for 400 epochs with a total batch size of 64 and learning rate (lr) 1e51e-5 with linear warmup and step reduce lr scheduler. As for the query denoising settings, the noise rate for the curve type and parameters is $0.2$ and $0.3$, respectively, with 5 denoising groups. The transformer encoder and decoder contain 6 and 8 transformer layers, respectively, with eight attention heads and latent model dimension 256.
Evaluation Metrics. We adopt command type accuracy ($Acc^{SKH}_{t}$) and parameter accuracy ($Acc^{SKH}_{p}$) for quantitative evaluations of the sketch prediction following (GONZALEZLLUCH2021103460). For networks that represent sketch by the implicit nueral field, the tools for convert sketch SDF to command are developed based on the tool by SECAD (Li_Guo_Zhang_Yan_2023). Furthermore, following Chen_Zhang_2019, we also report chamfer distance ($CD$), edge chamfer distance ($ECD$), normal consistency ($NC$), and the number of generated primitives ($\Delta\#P$) to measure the quality of the recovered 3D geometry, more details of the quantitative metrics are shown in the Appendix B.
@@ 4.2 Main Results
[b] DeepCAD Fusion 360 Gallery Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD ECDECD NCNC IR $\Delta P$ $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD ECDECD NCNC IR $\Delta P$ Methods of Representing Sketch with Implicit SDF Encoding ExtrudeNet 34.58% 31.71% 0.379 0.962 0.849 24.19% 34.34 37.81% 34.39% 0.671 0.808 0.809 23.83% 27.47 Point2Cyl 41.37% 39.41% 0.489 1.027 0.819 3.91% 26.18 42.98% 41.16% 0.529 0.769 0.769 3.87% 26.96 SECAD-Net 45.91% 41.37% 0.341 0.868 0.861 8.03% 32.78 46.82% 42.97% 0.449 0.684 0.813 7.82% 28.61 Generation Methods Based on Language Model DeepCAD 82.61% 73.36% 0.898 1.883 0.823 14.12% 5.89 77.31% 69.81% 7.128 8.729 0.719 13.18% 7.47 HNC-CAD 84.31% 76.71% 0.827 1.064 0.846 6.01% 4.43 80.62% 72.19% 4.381 5.571 0.748 5.92% 6.14 Methods Based on Primitive Fitting UCSG-Net - - 1.849 1.174 0.820 - 14.19 - - 0.952 1.277 0.770 - 11.17 CSG-Stump - - 3.031 0.755 0.828 - 19.46 - - 0.781 0.991 0.744 - 13.08 Ours 96.14% 86.81% 0.312 0.581 0.897 3.71% 4.14 94.17% 83.52% 0.392 0.571 0.839 3.62% 5.17
@@
To demonstrate its effectiveness, we compare the proposed model to multiple existing methods, covering the methods of representing sketch with SDF encoding (Li_Guo_Zhang_Yan_2023; Ren_Zheng_Cai_Li_Zhang_2022; 9880393), the generation methods based on language model (Wu_Xiao_Zheng_2021; xu2023hierarchicalneuralcodingcontrollable), and methods based on primitive fitting (kania2020ucsgnetunsuperviseddiscovering; ren2021csgstumplearningfriendlycsglike). The direct reconstruction results are visualized for comparison.
Table 1 summarizes the main results, and we can find that sketch curve type and parameter accuracy of the Point2Primitive  are much higher than the other methods while the CD metric of the Point2Primitive  is lower. The smallest CD and ECD value indicates that the presented CAD reconstruction method achieves better geometry fidelity while preserving accurate and sharp shape edges. This shows that by directly recovering curves of the extrusion primitive, our method can achieve better performance than the other extrusion-segmentation methods that represent the sketch with an implicit field. As for the number of generated primitives, our method is closer to human designs (The smallest $\Delta P$ value). This improvement is due to the explicit fine-grained reconstruction from each extrusion primitive; every parameter of the curves and extrusion operations are all predicted and optimized.
@@ 4.3 Visualization Results
Figure 3: Visual comparison between reconstruction results on the DeepCAD (the first 4 lines) and Fusion360 gallery dataset (the last 4 lines).
Methods $Acc^{SKH}_{t}$ $Acc^{SKH}_{p}$ CDCD ECDECD NCNC $\Delta P$
Methods of Representing Sketch with Implicit SDF Encoding
ExtrudeNet 28.39% 25.61% 0.614 1.117 0.776 36.14
Point2Cyl 34.47% 30.25% 0.518 1.065 0.791 27.96
SECAD-Net 36.61% 31.43% 0.437 1.079 0.806 34.18
Generation Method Based on Language Model
DeepCAD 61.41% 43.71% 5.919 6.883 0.708 7.16
HNC-CAD 63.39% 53.29% 6.864 7.064 0.711 6.56
Methods Based on Primitive Fitting
UCSG-Net - - 2.146 1.273 0.797 14.81
CSG-Stump - - 3.681 0.958 0.804 20.16
Ours 94.67% 83.81% 0.410 0.607 0.819 4.97
Table 2: Evaluation results on the Augmented DeepCAD dataset.
Figure 3 demonstrates the visualized reconstruction results on the DeepCAD and Fusion 360 Gallery datasets, respectively. It can be seen that the results of our method are more compact and complete, while the edges are sharp and sketches are accurate. This shows that our method can produce CAD reconstruction of high geometrical fidelity. The methods based on the language model can also produce accurate geometry structure, but the error of the parameters in the sequence results in some models of poor compactness. The primitive detection methods (USCG and Stump) fail to predict some holes in the model. Still, the edges are critically more accurate than the ExtrudeNet, whose reconstructions are less complete. Please refer to Appendix I for more visualizations.
@@ 4.4 Comparison on Augmented DeepCAD Dataset
Metrics Ours Point2Cyl
w/o $\mathcal{L}_{\text{skh}}$ w/ w/o $\mathcal{L}_{\text{skh}}$ w/
Ext. Seg IoU 0.814 0.862 0.801 0.817
BB. Seg Acc 0.902 0.902 0.867 0.867
EA. Angle Err $\downarrow$ 7.416 7.173 8.391 8.267
Fit Ext. $\downarrow$ 0.0771 0.0527 0.0791 0.0741
(a) Quantitative comparison on DeepCAD.
(b) There constructed sketches by Point2Cyl and Point2Primitive.
(c) The extrusion segmentation by Point2Cyl and Point2Primitive.
Figure 5: Comparison with Point2Cyl.
Table 2 demonstrates the quantitative results of the augmented DeepCAD dataset. We add the original shape (Origin) to the last column in the visualized results as shown in Figure 4. It can be seen that the methods based on the language model see a direct drop in the metrics of the sketch command type and command parameter, leading to a significant increase in the CD metric. Some reconstruction results by the language model preserve similarities to the original shape, leading to false reconstruction. Please refer to Appendix I for more visualizations.
The generated results are similar to the ground truth but with low geometry accuracies. On the other hand, reconstruction methods can still produce results that satisfy geometry fidelity. This implies that the methods based on the language model are less generalizable than the other methods. Also, both the quantitative metrics and the visual results verify that the proposed method can be generalized to reconstruction tasks beyond the train-test dataset.
@@ 4.5 Detailed Comparison with Point2Cyl
In Table 5(a), we present the quantitative comparison between our method and the most related approach, Point2Cyl. Compared with Point2Cyl, the metrics of the extrusion (Ext.) and barrel/based (BB.) segmentation are improved by $5.6\%$ and $3.5\%$, respectively, leading to $13\%$ and $32\%$ reduction on the extrusion axis (EA.) angle and fitting (Fit Ext.) error, respectively. This shows that by supervising the extrusion segmentation with curve loss (type and parameter loss), more geometrical information is introduced, leading to better geometry accuracy.
We visualized the sketch predicted by Point2Cyl and our method, as shown in Figure 5(b). It can be seen that because Point2Cyl represents the sketch by SDF, the edges of the sketch profiles are curved, while the counterpart of Point2Primitive  is sharp and accurate. In addition, we present more visualizations of the predicted sketch by Point2Primitive  in the Appendix C. Figure 5(c) (b) demonstrates the segmentation results by the two methods. It can be seen that even though the multiple parts of one extrusion are labeled with different extrusion labels, the segmentation network can still learn the geometry representation and cluster the points as the same extrusion. This implies that our method can predict simplified primitives that are closer to human design.